OpenStudy (anonymous):

If I have log base 2 the square root of c is that the same as log base 2 c to the power of 1/2?

ganeshie8 (ganeshie8):

\(\large \log_2 \sqrt{c}\)

ganeshie8 (ganeshie8):

\(\large (\large \log_2 c)^\frac{1}{2} \)

ganeshie8 (ganeshie8):

your two expressions are like above ?

OpenStudy (anonymous):

the first is, I am asking to simplify does it become log2 c1/2 (meaning power of 1/2) if that makes sense

OpenStudy (anonymous):

\[\log_{2} c ^{1/2}\]

ganeshie8 (ganeshie8):

\(\large \log_2 \sqrt{c}\) \( = \large \log_2 (c)^\frac{1}{2}\) \( = \large \frac{1}{2} \log_2 (c)\)

ganeshie8 (ganeshie8):

yes you're right :)

OpenStudy (anonymous):

Ahhhhh thank you :)

ganeshie8 (ganeshie8):

np :) we're donw with that. but as you can see, below two expressions are not same :- \(\large \log_2 \sqrt{c}\) \(\large (\log_2 c)^\frac{1}{2} \)

ganeshie8 (ganeshie8):

i hope you see why they're not same

OpenStudy (anonymous):

No because then if would be squaring the \[\log_{2} c \] component yes?

ganeshie8 (ganeshie8):

perfect, you mean square rooting !

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

good :)

OpenStudy (anonymous):

Thank you again, I struggle a bit with logs sometimes, I am sure it will make sense one day LOL

ganeshie8 (ganeshie8):

logs are tricky in the beginning... but i can see you're pretty good in these :)

OpenStudy (anonymous):

hahaha thanks :)

ganeshie8 (ganeshie8):

you're wlcme !